Fck:

Given the equation f(x) = x^5 - 10x^4 + 35x^3 - 50x^2 + 24x - 3, find the x-intercepts and classify each as rational, irrational, or imaginary. Then, use calculus to find the intervals where the function is increasing or decreasing, as well as the points of inflection. Finally, use the Intermediate Value Theorem to determine whether or not the function has any real roots between certain intervals.

Fck:

@ultrilliam

Fck:

@vocaloid

Fck:

@shadow

Fck:

@thisgirlpretty

Fck:

@elsa213

Fck:

@thekaylahope

Fck:

@fernyfernyferny

Fck:

@alanaaaaaaa

GlamorousGoddess:

@k1ngofpadlet

K1NGofPadlet:

When finding the x-intercepts of the function f(x) = x^5 - 10x^4 + 35x^3 - 50x^2 + 24x - 3, set f(x) equipollent to zero and solve for x. This gives you: x^5 - 10x^4 + 35x^3 - 50x^2 + 24x - 3 = 0 Infelicitously, there is no facile way to solve this equation algebraically, so you then resort to numerical methods or graphing to approximate the roots. However, you can determine the possible rational roots of the function by utilizing the Rational Root Theorem. According to the theorem, the possible rational roots of the function are all factors of the constant term (in this case, 3) divided by all factors of the leading coefficient (in this case, 1). That gives a list of possible rational roots: ±1, ±3. However, it's paramount to note that the function could additionally have irrational or imaginary roots, which would be more arduous to find. Consequently, the x-intercepts could be relegated as rational, irrational, or imaginary, depending on the genuine roots of the function.

K1NGofPadlet:

Now, to find the intervals where the given function f(x) = x^5 - 10x^4 + 35x^3 - 50x^2 + 24x - 3 is incrementing or decrementing, find its derivative and determine its sign for different intervals. The derivative of the function is given by f'(x) = 5x^4 - 40x^3 + 105x^2 - 100x + 24. To find where the function is incrementing, find the intervals where the derivative is positive. f'(x) > 0 5x^4 - 40x^3 + 105x^2 - 100x + 24 > 0 Utilizing calculus, factorize f'(x) and find its roots: f'(x) = (x - 1)(5x^3 - 35x^2 + 70x - 24) The root of the first factor is x = 1. To find the roots of the second factor, utilize polynomial division or numerical methods. By utilizing numerical methods, you can find that the other three roots are approximately x = 1.349, x = 2.694, and x = 3.459. Consequently, the function is incrementing on the intervals (-∞, 1), (1.349, 2.694), and (3.459, ∞). To find where the function is decrementing, find the intervals where the derivative is negative. f'(x) < 0 5x^4 - 40x^3 + 105x^2 - 100x + 24 < 0 The intervals where the function is decrementing are (1, 1.349), (2.694, 3.459).

K1NGofPadlet:

For the points of inflection, find the second derivative of the function and then find where it equals zero or fails to subsist. So let's commence by finding the first derivative: f'(x) = 5x^4 - 40x^3 + 105x^2 - 100x + 24 Now let's find the second derivative: f''(x) = 20x^3 - 120x^2 + 210x - 100 Next, you require to find where the second derivative equals zero or fails to subsist: f''(x) = 0 20x^3 - 120x^2 + 210x - 100 = 0 you can factor out 5 to simplify the calculation: 5(4x^3 - 24x^2 + 42x - 20) = 0 Utilizing the rational root theorem, you can find that the possible rational roots are ±1, ±2, ±4, ±5, ±10, ±20. Checking them piecemeal, you'll find that x = 1 and x = 2 are roots of the equation. So you can factor the equation as: 5(x - 1)(x - 2)(4x - 5) = 0 thus the points of inflection are x = 1, x = 2, and x = 5/4.

K1NGofPadlet:

To utilize the Intermediate Value Theorem (IVT), you'll want to first identify two values of x, let's call them a and b, such that f(a) and f(b) have antithesis signs. This would implicatively insinuate that f(x) must cross the x-axis at some point between a and b, which would betoken there is at least one genuine root in that interval. Sob you can better understand this concept, let's take an optical canvassing of the function f(x) = x^3 - x - 1. By utilizing a graphing calculator, I can visually perceive that there is a genuine root between the intervals [1,2] and [-2,-1]. I can corroborate this by plugging in the values for a and b: f(1) = 1 - 1 - 1 = -1 f(2) = 8 - 8 + 12 - 6 + 1 = 7 f(-2) = -8 + 16 - 12 + 6 - 1 = -1 f(-1) = -1 + 1 - 1 = -1 Since we have antithesis signs for f(1) and f(2), and for f(-2) and f(-1), this tells us that there must be at least one authentic root between the intervals [1,2] and [-2,-1]. So, to apply this same method to the function f(x) = x^5 - 10x^4 + 35x^3 - 50x^2 + 24x - 3, we require to identify two values of x that will give us antithesis signs when we plug them in. We can commence by plugging in the values 0 and 1: f(0) = -3 f(1) = -3 Uh oh, it looks akin to the function has the same sign at x = 0 and x = 1. We have antithesis signs for f(-1) and f(2), which denotes there must be at least one genuine root between the intervals [-1,2]. We can perpetuate to utilize this method and endeavor plugging in some other values to identify other intervals with authentic roots.

Fck:

But thank you guys bro needed the help 😆

K1NGofPadlet:

@rizzismybizz wrote:
@k1ngofpadlet wrote:
When finding the x-intercepts of the function f(x) = x^5 - 10x^4 + 35x^3 - 50x^2 + 24x - 3, set f(x) equipollent to zero and solve for x. This gives you: x^5 - 10x^4 + 35x^3 - 50x^2 + 24x - 3 = 0 Infelicitously, there is no facile way to solve this equation algebraically, so you then resort to numerical methods or graphing to approximate the roots. However, you can determine the possible rational roots of the function by utilizing the Rational Root Theorem. According to the theorem, the possible rational roots of the function are all factors of the constant term (in this case, 3) divided by all factors of the leading coefficient (in this case, 1). That gives a list of possible rational roots: ±1, ±3. However, it's paramount to note that the function could additionally have irrational or imaginary roots, which would be more arduous to find. Consequently, the x-intercepts could be relegated as rational, irrational, or imaginary, depending on the genuine roots of the function.
Yea smart bro. This was a better shortcut. Mine was going to be an hour's worth of work because I don't know what his knowlegeability is.
No worries man. Have a good one.

K1NGofPadlet:

@fck wrote:
But thank you guys bro needed the help 😆
No problem my guy.

K1NGofPadlet:

@glamorousgoddess wrote:
@k1ngofpadlet
Thanks for the @ it's not often people ask questions anymore. 😄🙏